OpenStudy (anonymous):

The functions f(x) = 2x and g(x) = 2|x| are equal on the domain x > 0. A. True B. False

OpenStudy (anonymous):

true?

OpenStudy (anonymous):

wait no false because of the lxl

Parth (parthkohli):

Let's try it with -2 maybe. -2 * 2 = -4 2|-2| = 2 * 2 = 4 False!

OpenStudy (anonymous):

yea! :)